# A Simple Unified Uncertainty-Guided Framework for Offline-to-Online Reinforcement Learning

### v2

@@ 4 SUNG
In this section, we present a simple unified uncertainty-guided framework (SUNG) for offline-to-online RL, as depicted in Fig. 1. Concretely, SUNG actively quantifies the state-action uncertainty by estimating a density function with VAE. Then, under the guidance of uncertainty, SUNG incorporates an optimistic exploration strategy and an adaptive exploitation method for online finetuning. Combined with OORB, we finally arrive at the complete algorithm of SUNG.
4.1 Uncertainty Quantification with Density Estimation
@@
Many prior works typically utilize Q ensembles to qualify uncertainty Wu et al. (2021); Mark et al. (2022); Ciosek et al. (2019); Lee et al. (2021). However, the ensemble technique may significantly increase computational costs. Thus, in this work, we utilize a simple yet effective approach by adopting VAE Kingma and Welling (2013) as a state-action visitation density estimator for uncertainty quantification. Specifically, given an encoder $q_{\varphi}(z|s,a)$ and a decoder $p_{\psi}(s,a|z)$ with parameters $\varphi$ and $\psi$, respectively, we can optimize them with the evidence lower bound (ELBO)
$\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)=\mathbb{E}_{q_{\varphi}(z|s,a)}\left[-\log p_{\psi}(s,a|z)\right]+D_{\mathrm{KL}}\left[q_{\varphi}(z|s,a)||p(z|s,a)\right],$ (6)
@@
where $p(z|s,a)=\mathcal{N}(0,I)$ is a fixed prior. Then, we utilize the negative log likelihood of the state-action density for uncertainty quantification, i.e., $\mathcal{U}(s,a)\overset{\rm{def}}{=}-\log p(s,a)\approx\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$, as we know $\log p(s,a)=-\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)+D_{\mathrm{KL}}[q_{\varphi}(z|s,a)||p_{\psi}(z|s,a)]$. Intuitively, both the reconstruction loss and KL divergence in ELBO indicate the state-action uncertainty, i.e., whether a state-action pair can be well captured by the learned distribution.
Although it can be theoretically shown that a bias exists between ELBO and the negative log likelihood of the state-action density, a prior work has shown that such approximation is empirically valid Wu et al. (2022). Thus, we utilize ELBO for uncertainty quantification in practice, and leave the exploration of importance sampling techniques Kingma et al. (2019); Rezende et al. (2014) for further bias reduction as future work.
4.2 Optimistic Exploration via Bi-level Action Selection
@@
A main line of model-free offline RL methods perform conservative objectives to penalize OOD actions during offline pretraining. However, such conservative objectives inherently limit agents’ exploration power, which impedes agents to fully benefit from online finetuning Mark et al. (2022). To overcome this limitation, our goal is to properly measure the informativeness of unseen actions, and then select the most informative action for exploration. To achieve this, we propose an optimistic exploration strategy based on the principle of optimism in the face of uncertainty Brafman and Tennenholtz (2002); Audibert et al. (2009). The key insight is that an informative action is expected to possess both high Q value and high uncertainty.
@@
Existing works typically estimate epistemic uncertainty via Q ensembles and derive an upper confidence bound (UCB) to direct the exploration Chen et al. (2017); Ciosek et al. (2019); Lee et al. (2021). Different from them, we extend this idea by utilizing $\mathcal{U}(s,a)$ for uncertainty quantification. To avoid inaccurate estimates for Q value and uncertainty, we only select near-on-policy actions for exploration. Consequently, the objective of the exploration policy $\pi_{E}(s)$ is to maximize the informativeness while remaining near-on-policy, i.e.,
@@
$\mathrm{s.t.}\ ||a-\pi_{\phi}\left(s\right)||^{2}\leq\delta,$
where $\beta$ controls the optimism level and $\delta$ controls the on-policyness. As such, the derived behavior policy can not only increase the chance of executing informative actions according to the principal of optimism in the face of uncertainty, but also guarantee the on-policyness to ensure the stability.
However, such an optimization problem is impractical to solve due to two main challenges. On the one hand, it is not straightforward to find optimal action in high-dimensional continuous action space, which is impossible to enumerate exhaustively. On the other hand, it is hard to set a proper value for optimism level $\beta$ because Q value and uncertainty have different value ranges across different tasks, which necessitates task-specific hyper-parameter tuning.
To mitigate both challenges above, we propose a simple yet effective bi-level action selection mechanism. First, we generate a candidate action set $\mathcal{C}=\{a_{i}\}_{i=1}^{N}$ with $N$ candidate actions. Here, the action is generated by sampling $a_{i}\sim\pi_{\phi}(\cdot|s)$ for stochastic policy, while by adding sampled Gaussian noise $a_{i}=\pi_{\phi}(s)+\epsilon_{i},\epsilon_{i}\sim\mathcal{N}(0,\delta)$ for deterministic policy. Then, we rank the candidate actions according to the Q-values (or uncertainty) to select top-$k$ candidate actions as a finalist action set $\mathcal{C}_{f}$. Finally, we construct a categorical distribution according to the uncertainty (or Q values) to select the action from $\mathcal{C}_{f}$ for interacting with the environment:
$\mathrm{P}(a_{i}):=\frac{\mathrm{exp}\left(\mathcal{U}(s,a_{i})/\alpha\right)}{\sum_{j}\mathrm{exp}\left(\mathcal{U}(s,a_{j})/\alpha\right)},\forall i\in[1,...,k],$ (8)
where $\alpha$ is the softmax temperature. By altering ranking criteria for the finalist action set and $k$, we can flexibly adjust the preference for high Q value or high uncertainty to achieve the desired trade-off. In practice, we observe that policy constraint based offline RL methods can accommodate preference for either high Q value or high uncertainty. However, value regularization based methods may diverge when preference for Q value is present. We detailedly discuss this phenomenon in Appendix B.
@@
4.3 Adaptive Exploitation with OOD Sample Identification
Action distribution shift poses a significant challenge for offline RL algorithms Fujimoto et al. (2019); Kumar et al. (2019); Levine et al. (2020); Kumar et al. (2020), as the bootstrapping term in policy evaluation involves actions derived from the learned policy $\pi_{\phi}$. This can result in the extrapolation error due to querying Q functions with OOD actions, leading to biased policy improvement towards these OOD actions with erroneously high Q values. Note that offline RL methods do not suffer from state distribution shift during training, since policy evaluation only queries Q functions with states present in the offline dataset.
However, during online finetuning, both state and action distribution shifts occur since agents may encounter unfamiliar state-action regimes that fall outside of the support of the dataset. Worse still, as the proposed optimistic exploration strategy follows the principle of optimism in the face of uncertainty, the state-action distribution shift may be further exacerbated. Preliminary experiments from Lee et al. (2022); Mark et al. (2022); Zheng et al. (2023) show that this may erase the good initialization obtained from offline pretraining.
To tackle the state-action distribution shift issue, we propose an adaptive exploitation method with OOD sample identification. The underlying idea is developed from the observations in previous works Lee et al. (2022); Mark et al. (2022); Zheng et al. (2023) that finetuning with continual offline RL objectives typically derives stable but limited performance, while finetuning with online RL objectives typically derives unstable performance due to the distribution shift caused by OOD state-action pairs. Therefore, we propose a novel approach that leverages the benefits of both objectives by utilizing conservative offline RL objectives for state-action pairs with high uncertainty while aggressive online RL objectives for state-action pairs with low uncertainty. As shown in Eq. (5), offline RL objectives consist of a standard online RL objective $\mathcal{L}_{\mathcal{F}}(\Theta)$ and a regularizer $\mathcal{R}_{\mathcal{F}}(\Theta)$. Thus, we can derive the objectives for adaptive exploitation by introducing an uncertainty-guided OOD sample identifier $\mathcal{I}(s,a)$:
@@ 4.3 Adaptive Exploitation with OOD Sample Identification
$\mathcal{I}(s,a):=\left\{\begin{array}[]{ll}1,&\mathrm{if}(s,a)\in\mathcal{D}_{\mathrm{OOD}},\\ 0,&\mathrm{else}.\\ \end{array}\right.$ (10)
By setting $p=0$, the training objective is precisely the offline RL objective, while setting $p=100$ results in the online RL objective. Thus, by tuning the value of $p$, the proposed adaptive exploitation method can attain the desired trade-off between performance and stability.
@@
13:     Sample the action $a_{t}\sim\mathrm{P}(a_{i}),a_{i}\in\mathcal{C}_{f}$ in Eq.(8) according to the uncertainty (or Q value)
14:     Execute the action $a_{t}$, collect $s_{t+1}$ and ${r_{t}}$ from the environment
@@
4.4 Algorithm Implementation Details
@@
Offline-to-Online Replay Buffer. Previous studies Lee et al. (2022); Zheng et al. (2023) have highlighted the significance of incorporating both offline and online data during finetuning. Here, we adopt the offline-to-online replay buffer proposed in Zheng et al. (2023) due to its simplicity. Specifically, OORB consists of two replay buffers: a fully online buffer that collects transitions from online interactions, and an offline buffer that stores transitions from both offline dataset and online interactions. During training, we randomly sample transitions from OORB with a Bernoulli distribution, where the probability of sampling from the fully online buffer is $p_{\mathrm{OORB}}$, and the probability of sampling from the offline buffer is $1-p_{\mathrm{OORB}}$.
@@
SUNG Framework for Offline-to-Online RL. Putting everything together, we summarize the full framework in Algorithm 1. SUNG first pretrains agents and VAE using offline dataset. Then, it turns to online finetuning with the proposed optimistic exploration strategy and adaptive exploitation method. We also provide two implementation cases of bridging TD3+BC and TD3, and bridging CQL and SAC in Appendix A.
5 Experiments
@@
In this section, we provide empirical results to verify the effectiveness of our framework by answering three primary questions: (1) How does SUNG compare with other state-of-the-arts when combined with different offline RL methods, across various types of tasks and datasets? (2) How do the key components of SUNG influence the overall offline-to-online performance? (3) How should we set hyper-parameters for SUNG in practice, and how challenging is this task?
5.1 Experimental Setup
@@
Settings. We evaluate on the D4RL benchmark Fu et al. (2020), which provides various continuous-control tasks and datasets. We focus on MuJoCo and AntMaze domains. We first perform 1M gradient steps for offline pretraining, and then perform 100K environment steps for online finetuning. While some prior works Kostrikov et al. (2022); Zhang et al. (2023); Zheng et al. (2023) take 1M environment steps for finetuning, we argue that 1M environment steps are even enough for an online RL agent to achieve expert-level performance. Thus, we believe finetuning for 100K environment steps is a more reasonable setting, which is also adopted by Lyu et al. (2022); Zheng et al. (2023).
@@
Backbone Offline RL Methods. To evaluate the generality of SUNG, we choose two representative offline RL methods as the backbone for pretraining, i.e., TD3+BC Fujimoto and Gu (2021) and CQL Kumar et al. (2020). Concretely, TD3+BC extends TD3 with behavior cloning based policy constraint, while CQL extends SAC with value regularization. For AntMaze domains, we substitute TD3+BC with SPOT Wu et al. (2022) due to its inferior performance. SPOT is another extension of TD3 but with density guided policy constraint.
Baselines. We compare SPOT with the following baselines: (1) offline-ft leverages online interactions by performing offline RL objectives. (2) online-ft leverages online interactions by performing online RL objectives that correspond to the ones used for pretraining. (3) BR Lee et al. (2022) prioritizes near-on-policy transitions from the replay buffer. (4) O3F Mark et al. (2022) utilizes knowledge contained in value functions to guide exploration. (5) APL Zheng et al. (2023) performs adaptive policy learning by incorporating different advantages of offline and online data. (6) PEX Zhang et al. (2023) trains a new policy from scratch by utilizing the pre-trained policy for exploration. Note that we also include ODT Zheng et al. (2022) for comparison with a seperate subsection in Appendix D.2. See Appendix C for further implementation details.
5.2 Comparisons on D4RL Benchmarks

### v3

@@ IV SUNG
In this section, we present a simple unified uncertainty-guided framework (SUNG) for offline-to-online RL, as depicted in Fig. 1. Concretely, SUNG quantifies the state-action uncertainty by estimating a density function with VAE. Then, under the guidance of uncertainty, SUNG incorporates an optimistic exploration strategy and an adaptive exploitation method for online finetuning. Finally, we summarize the full framework in Algorithm 1.
IV-A Uncertainty Quantification with Density Estimation
@@
Many prior works typically utilize Q ensembles to qualify uncertainty [16, 11, 53, 13]. However, the ensemble technique may significantly increase computational costs. Thus, in this work, we utilize a simple yet effective approach by adopting VAE [17] as a state-action visitation density estimator for uncertainty quantification. Specifically, VAE consists of an encoder $q_{\varphi}(z|s,a)$ and a decoder $p_{\psi}(s,a|z)$ with parameters $\varphi$ and $\psi$, respectively. The VAE is optimized with the evidence lower bound (ELBO):
$\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$ $=\mathbb{E}_{q_{\varphi}(z|s,a)}\left[-\log p_{\psi}(s,a|z)\right]$ (7)
$+D_{\mathrm{KL}}\left[q_{\varphi}(z|s,a)||p(z)\right],$
where $p(z)=\mathcal{N}(0,I)$ is a fixed prior distribution. In ELBO, the first term is the reconstruction loss, while the second term is the KL divergence that regularizes the posterior distribution $q_{\varphi}(z|s,a)$ to be close to the prior $p(z)$.
After training a VAE based on the offline dataset $\mathcal{D}$, we utilize the negative log likelihood of the state-action density for uncertainty quantification, which can be further approximated by ELBO, i.e.,
$\mathcal{U}(s,a)$ $\overset{\rm{def}}{=}-\log p(s,a)$ (8)
$=\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)-D_{\mathrm{KL}}[q_{\varphi}(z|s,a)||p_{\psi}(z|s,a)]$
$\approx\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi).$
@@
Intuitively, both the reconstruction loss and KL divergence in ELBO indicate the uncertainty for state-action space, i.e., whether a state-action pair can be well captured by the learned distribution in VAE. We provide more justification of such approximation in Appendix B in the supplementary material.
Note that VAEs have been widely adopted in previous offline RL works—such as BCQ [9], SPOT [28], and CPQ [46]—to model the behavior policy $\pi_{\beta}$ of the offline dataset $\mathcal{D}$. This modeled policy is then used to constrain the learned policy from offline RL, preventing it from deviating from the offline policy distribution too much. Different from them, we adopt VAE to estimate the visitation density of the state-action pair as uncertainty quantification, which can guide optimistic exploration and adaptive exploitation in offline-to-online RL. We present a illustrative comparison of them in the supplementary material.
IV-B Optimistic Exploration via Bi-level Action Selection
@@
A main line of model-free offline RL methods perform conservative objectives to penalize OOD actions during offline pretraining. However, such conservative objectives inherently limit agents’ exploration power, which impedes agents to fully benefit from online finetuning [11]. To overcome this limitation, our goal is to properly measure the informativeness of unseen actions, and then select the most informative action for exploration. To achieve this, we propose an optimistic exploration strategy via bi-level action selection mechanism. In this subsection, we first formulate the objective of the optimistic exploration strategy, and then propose a practical bi-level action selection mechanism to find approximate solutions of this optimization objective.
To overcome the constrained exploratory behavior issue, we propose an optimistic exploration strategy based on the principle of optimism in the face of uncertainty, which is theoretically grounded on the upper confidence bound (UCB) algorithm [18, 19] in the bandit setting. The key insight here is that we encourage the agents to select informative actions with both high Q value and high uncertainty.
@@
Previous works typically estimate epistemic uncertainty via Q ensembles and utilize the UCB algorithm to direct the exploration [54, 53, 13]. Different from them, we extend this idea by utilizing $\mathcal{U}(s,a)$ for uncertainty quantification. To avoid inaccurate estimates for Q value and uncertainty, we only select near-on-policy actions for exploration. Consequently, the objective of the exploration policy $\pi_{E}(s)$ is to maximize the informativeness while remaining near-on-policy, i.e.,
@@
[…]
where $\beta$ controls the optimism level and $\delta$ controls the on-policyness, given a general assumption of smoothness in the physical transition model [49]. As such, the derived behavior policy can not only encourage informative actions according to the principal of optimism in the face of uncertainty, but also guarantee the on-policyness to ensure the stability.
However, such an optimization problem is impractical to solve due to two main challenges: (1) It is not straightforward to find optimal action in high-dimensional continuous action space, which is impossible to enumerate exhaustively. (2) It is hard to set a proper value for optimism level $\beta$ because Q value and uncertainty have different value ranges across different tasks, which necessitates task-specific hyper-parameter tuning.
To mitigate both challenges above, we propose a practical bi-level action selection mechanism. The key idea here is to approximate the intractable constrained optimization by breaking it into two simpler stages: first narrowing down a promising subset of near-on-policy candidate actions using one criterion (Q or uncertainty), and then probabilistically selecting among them using the complementary criterion.
Specifically, we first generate a candidate action set $\mathcal{C}=\{a_{i}\}_{i=1}^{N}$ with $N$ candidate actions. Here, the action is generated by sampling $a_{i}\sim\pi_{\phi}(\cdot|s)$ for stochastic policy, while by adding sampled Gaussian noise $a_{i}=\pi_{\phi}(s)+\epsilon_{i},\epsilon_{i}\sim\mathcal{N}(0,\delta)$ for deterministic policy. Then, we rank the candidate actions according to the Q-values (or uncertainty) to select top-$k$ candidate actions as a finalist action set $\mathcal{C}_{f}$. Finally, we construct a categorical distribution according to the uncertainty (or Q values) to select the action from $\mathcal{C}_{f}$ for interacting with the environment:
$\mathbb{P}(a_{i}|s):=\frac{\mathrm{exp}\left(\mathcal{U}(s,a_{i})/\alpha\right)}{\sum_{j}\mathrm{exp}\left(\mathcal{U}(s,a_{j})/\alpha\right)},\forall i\in[1,...,k],$ (10)
where $\alpha$ is the softmax temperature. Intuitively, when we set the ranking criteria for the finalist action set as Q values, if we set $k=1$, the exploration policy would greedily choose high-value actions; if we set $k=N$, the exploration policy would favor those with high uncertainty actions. As such, by altering ranking criteria for the finalist action set and $k$, we can flexibly adjust the preference for high Q value or high uncertainty to achieve the desired trade-off.
As discussed in [11], value regularization based methods may diverge when preference for Q value is present. To address this, we establish the ranking criteria for the finalist action set as uncertainty for value regularization-based methods and as Q value for other offline RL methods.
@@
IV-C Adaptive Exploitation with OOD Sample Identification
Action distribution shift poses a significant challenge for offline RL algorithms [9, 26, 1, 10], as the bootstrapping term in policy evaluation involves actions derived from the learned policy $\pi_{\phi}$. This can result in the extrapolation error due to querying Q functions with OOD actions, leading to biased policy improvement towards these OOD actions with erroneously high Q values. Note that model-free offline RL methods do not suffer from state distribution shift during training, since policy evaluation only queries Q functions with states present in the offline dataset.
However, during online finetuning, both state and action distribution shifts occur since agents may encounter unfamiliar state-action regimes that fall outside of the support of the dataset. Worse still, as the proposed optimistic exploration strategy follows the principle of optimism in the face of uncertainty, the state-action distribution shift may be further exacerbated. Preliminary experiments from [6, 11, 12] show that this may erase the good initialization obtained from offline pretraining.
To tackle the state-action distribution shift issue, we propose an adaptive exploitation method with OOD sample identification. The underlying idea is developed from the observations in previous works [6, 11, 12] that finetuning with continual offline RL objectives typically derives stable but limited performance, while finetuning with online RL objectives typically derives unstable performance due to the distribution shift caused by OOD state-action pairs. Therefore, we propose a novel approach that leverages the benefits of both objectives by utilizing conservative offline RL objectives for state-action pairs with high uncertainty while aggressive online RL objectives for state-action pairs with low uncertainty. As shown in Eq. (6), offline RL objectives consist of a standard online RL objective $\mathcal{L}_{\mathcal{F}}(\Theta)$ and a regularizer $\mathcal{R}_{\mathcal{F}}(\Theta)$. Thus, we can derive the objectives for adaptive exploitation by introducing an uncertainty-guided OOD sample identifier $\mathcal{I}(s,a)$:
@@ IV-C Adaptive Exploitation with OOD Sample Identification
[…]
By setting $p=0$, the training objective is precisely the online RL objective, while setting $p=100$ results in the offline RL objective. Thus, by tuning the value of $p$, the proposed adaptive exploitation method can attain the desired trade-off between performance and stability.
A related work, CPQ [46], also employs a VAE to detect OOD actions and modifies the Bellman updates in Q-learning to mitigate overestimation for such actions. In contrast, SUNG leverages OOD sample identification to adaptively introduce regularization terms for OOD samples, aiming to balance performance and stability during online finetuning.
@@
13:  Sample the action $a_{t}\sim\mathbb{P}(a_{i}|s_{t}),a_{i}\in\mathcal{C}_{f}$ in Eq. (10) according to the uncertainty (or Q value)
14:  Execute the action $a_{t}$, collect next state $s_{t+1}$ and reward $r_{t}$} from the environment
@@
IV-D Algorithm Implementation Details
IV-D1 Offline-to-Online Replay Buffer
@@
Previous studies [6, 12] have highlighted the significance of incorporating both offline and online data during finetuning. Here, we adopt the offline-to-online replay buffer (OORB) proposed in [12] due to its simplicity. Specifically, OORB consists of two replay buffers: a fully online buffer that collects transitions from online interactions, and an offline buffer that stores transitions from both offline dataset and online interactions. During training, each transition in a minibatch is sampled using a Bernoulli-based source selection strategy: for each transition, we flip a Bernoulli coin with success probability $p_{\text{OORB}}$. If the coin lands heads (i.e., success), we sample a transition uniformly from the online buffer; otherwise, we sample uniformly from the offline buffer. This sampling procedure is repeated independently for every transition in the minibatch.
IV-D2 SUNG Framework for Offline-to-Online RL
@@
Summarizing the components outlined above, we present the complete framework in Algorithm 1. SUNG first pretrains agents and VAE estimator using the offline dataset. Then, it turns to online finetuning with the proposed optimistic exploration strategy and adaptive exploitation method.
Figure 2: D4RL environment. Above: MuJoCo tasks of halfcheetah, hopper and walker2d. Below: AntMaze tasks with u-shape, medium, large maze.
V Experiments
In this section, we provide empirical results to verify the effectiveness of our framework, aiming to answer the following four primary research questions (RQ):
@@ V Experiments
RQ1. How does SUNG compare with other state-of-the-arts when combined with different offline RL methods, across various types of tasks and datasets?
RQ2. How do the key components of SUNG influence the overall offline-to-online performance?
RQ3. How should we set hyper-parameters for SUNG in practice, and how challenging is this task?
RQ4. Can SUNG seamlessly integrate with other advanced RL techniques?
V-A Experimental Setup
V-A1 Environment and Dataset
We evaluate on the D4RL benchmark [20], which provides various continuous-control tasks and datasets. We focus on MuJoCo and AntMaze domains, as shown in Fig 2. For MuJoCo tasks, we follow prior work [7, 36] to omit the evaluation on medium-expert and expert datasets, where offline pre-trained agents can already achieve expert-level performance and do not need further finetuning. Thus, we only focus on three kinds of quality: random (r), medium (m), and medium-replay (m-r). For AntMaze tasks, we select all the six settings, including u-shape, medium and large maze. For fair comparison, all the evaluations use datasets of the latest released ”-v2” version that has fixed some potential bugs in the offline dataset.
V-A2 Evaluation Protocol
@@
We first perform 1M gradient steps for offline pretraining, and then perform 100K environment steps for online finetuning. While some prior works [32, 7, 12] take 1M environment steps for finetuning, we argue that 1M environment steps are even enough for an online RL agent to achieve expert-level performance. Thus, we believe finetuning for 100K environment steps is a more reasonable setting, which is also adopted by [29, 12]. All the experiments are repeated with 5 different random seeds. For each evaluation, we run 10 episodes for MuJoCo tasks and 100 episodes for AntMaze tasks. We employ the normalized score, i.e., D4RL score, as a metric to assess the agent’s performance. In this context, a score of 0 signifies random performance, while a score of 100 indicates expert-level performance. We report the average D4RL score over last 3 evaluations for MuJoCo tasks, and last 10 evaluations for AntMaze tasks.
V-A3 Backbone Offline RL Methods
@@
To evaluate the generality of SUNG, we choose two representative offline RL methods as the backbone for pretraining, i.e., TD3+BC [27] and CQL [10]. Concretely, TD3+BC extends TD3 with behavior cloning based policy constraint, while CQL extends SAC with value regularization. For AntMaze domains, we substitute TD3+BC with SPOT [28] due to its inferior performance. SPOT is another extension of TD3 but with a VAE-based density-guided policy constraint.
TABLE I: Hyper-parameters of online finetuning in SUNG.
TD3+BC CQL CQL-10
SUNG Candidate action num NN 100 100 20
Finalist action num $k$ 10 20 4
Softmax temperature $\alpha$ 1.0
OOD sample percentage pp 5 10 5
OORB Online buffer size 2e4
Offline buffer size 2e6
OORB probability $p_{\mathrm{OORB}}$ 0.1
V-A4 Implementation details and hyper-parameters
For VAE training, we follow the official implementation of BCQ [9] and SPOT [28]. Unlike these methods, which use VAEs to estimate the behavior policy’s action distribution, SUNG employs a VAE to model the state-action visitation density. Thus, we set the latent dimension of VAE to be twice the sum of the state and action dimensions. Note that we simply take the hyper-parameters reported in SPOT [28] and omit hyper-parameter tuning for VAE training. Accordingly, further performance improvement is expected from careful tuning.
For online finetuning, we follow the same standard architecture and the same hyper-parameter of TD3+BC, CQL, and SPOT for all the baselines. Note that we do not tune any hyper-parameter of backbone RL and architecture in SUNG for fair comparison. We present the detailed hyper-parameter setting of SUNG in Table I.
In terms of baseline reproduction, we strictly follow their official implementation and hyper-parameters reported in their original paper to finetune agents pre-trained with TD3+BC, CQL and SPOT for MuJoCo and AntMaze domains. One exception is that the heavy ensemble technique is removed in BR for fair comparison. For fair comparison, we compare SUNG with the full implementation of [6] as BRPQ in Section V-F. Note that we also equip the baseline with OORB to achieved improved performance for fair comparison.